# MeteoRA: Multiple-tasks Embedded LoRA for Large Language Models

### v1

@@ 4.2 Main results
Figures 3(a) and 3(b) demonstrate the performance of the MeteoRA models based on LlaMA2-13B and LlaMA3-8B, respectively, across the selected 28 tasks. We compare these models against three benchmarks for a comprehensive evaluation: the LLM loaded 28 LoRA adapters fine-tuned on each individual task via Huggingface PEFT as a reference model (PEFT), the model with LoRA fine-tuned using a mixed dataset aggregated from all tasks (LoRA-F), and the model trained with a balanced dataset (LoRA-B).
@@
| PEFT (reference) | 0.762 | 38.55 | 0.339 | 0.162 | 0.315 |
| MeteoRA (top-1) | 0.748 | 41.74 | 0.336 | 0.160 | 0.314 |
| MeteoRA (top-2) | 0.754 | 36.73 | 0.336 | 0.160 | 0.313 |
| LoRA-F | 0.729 | 41.27 | 0.318 | 0.136 | 0.294 |
| LoRA-B | 0.665 | 37.98 | 0.318 | 0.128 | 0.288 |
| Model | Accuracy | BLEU | ROUGE-1 | ROUGE-2 | ROUGE-L |
| PEFT (reference) | 0.817 | 45.32 | 0.341 | 0.164 | 0.317 |
| MeteoRA (top-1) | 0.806 | 44.98 | 0.337 | 0.158 | 0.313 |
| MeteoRA (top-2) | 0.811 | 45.64 | 0.338 | 0.158 | 0.314 |
| LoRA-F | 0.767 | 42.93 | 0.327 | 0.157 | 0.306 |
| LoRA-B | 0.750 | 38.47 | 0.343 | 0.171 | 0.320 |
The evaluation results indicate that, regardless of the base LLM, the MeteoRA models utilizing the top-1 strategy achieve performance very close to the reference model PEFT, while no explicit LoRA activation/deactivation is required in MeteoRA. Although LLMs with both LoRA-F and LoRA-B reach comparable performance on several certain tasks, they exhibit significantly poorer outcomes on others, such as logical_deduction. Additionally, MeteoRA employing the top-2 strategy, despite occasionally showing greater capability loss compared to MeteoRA with top-1 strategy, occasionally outperforms PEFT with adapters trained directly on the individual tasks. This suggests that the component in the loss function (Equation 4) becomes influential in these cases, indicating a beneficial mix of LoRA adapters from various tasks for future study.
@@ 4.3 Composte-n tasks
The evaluation results for these three tasks are illustrated in Table 3. Notice that only the MeteoRA with the top-2 strategy and LoRA with a balanced dataset effectively address these composite-n tasks. Subsequent discussions will therefore focus exclusively on these two models. Given that the format of our evaluation composite-n datasets resembles an exam paper, we evaluate the two models by following metrics for assessment: average number of questions attempted, average number of multiple-choice questions answered correctly, and average BLEU, ROUGE scores for non-multiple-choice questions. Observations from the composite-3 tasks reveal that although the MeteoRA model attempts slightly fewer questions than LoRA-B, it correctly answers a higher number of multiple-choice questions and achieves superior BLEU and ROUGE scores. As the task complexity increases in the composite-5 and composite-10 datasets, MeteoRA model outperforms LoRA-B in almost all metrics.
@@
To further validate the functionality of the Gating Network in the MeteoRA block, we stored the selection patterns of the LoRA adapters during the generating process of a composite-3 sample. The results are displayed in Figure 4. It is evident that the Gating Network, using the top-2 strategy, appropriately assigned greater weight to the corresponding LoRA adapters for the majority of the tokens, no matter in input or output. Notably, at the junctures where different tasks adjoin, the Gating Network correctly performed switching actions of LoRA adapters. This indicates that the Gating Network has indeed connected the tokens and the corresponding LoRA adapters.
To further validate the functionality of the Gating Network within the MeteoRA block, we analyze the selection patterns of the LoRA adapters during the generation process of a composite-3 sample. The results are depicted in Figure 4. It is evident that the Gating Network, utilizing the top-2 strategy, consistently assigned greater weight to the appropriate LoRA adapters for the majority of tokens, both in input and output phases. Specifically, during transitions between the two adjacent tasks, the Gating Network effectively switches LoRA adapters. This observation supports that the Gating Network successfully correlates tokens with their corresponding LoRA adapters, underscoring its effectiveness in managing dynamic task requirements.
@@ 5 Related Work
Multi-task fusion. Our proposed method falls into the field of LoRA adapter composition for multi-task fusion.
@@
The first category focuses on fusing the entire models. Researchers mainly study model ensembling and multi-task learning to achieve this goal. Existing works integrate the models under the setting of shared model architecture (Matena and Raffel, 2022; Jin et al., 2022; Wu et al., 2023b; Yadav et al., 2024). Others focus on merging models with various architectures or from different tasks. Both methods (Stoica et al., 2023; Liu et al., 2022) try to merge models that are trained for various tasks without additional training. The second category is more concerned with fusion in terms of the tasks. Ilharco et al. (2022) proposes a model editing method via task vectors. Sun et al. (2022) leverage in-context learning with few-shots to enhance the performance of unseen tasks. However, these methods require multi-task training or prior knowledge for the evaluation tasks. Our method could fuse off-the-shelf LoRA adapters with a lightweight fine-tuning for the Gate network in the MeteoRA module. During inference, none of the examples (zero-shot) are required for all individual tasks.
@@
Fusion under MoE. In the context of pretrain-fine-tune paradigm, PEFT becomes a common sense for developing Transformer-based LLM downstream applications. Directly fine-tuning on a fused dataset from various tasks is unable to achieve better performance Ling et al. (2024). Some works focus on leveraging existing LoRA adapters as off-the-shelf components, integrating them directly into a base LLM. For example, PEFT (Mangrulkar et al., 2022) and S-LoRA (Sheng et al., 2023) are frameworks aiming to fuse multiple LoRA adapters to one LLM. However, requiring explicit activation/deactivition during usage. Others (Huang et al., 2023; Yang et al., 2024; Feng et al., 2024; Chen et al., 2024; Wu et al., 2023c) propose LoRA fusion based on the concept of Mixture-of-experts the enhance the model’s ability for cross-domain tasks. However, the methods mainly focus on fusing LoRA adapters to the FFN module or Q in the attention module. Our method could fuse LoRA adapters that are applied to all types of linear layers in Transformer architecture. By leveraging the MoE architecture, the LLM’s capacity could be boosted with dynamic and timely LoRA switching during generation, especially for sequentially solving composite tasks.
@@ 6 Conclusions
This paper presents a multi-task embedded LoRA framework MeteoRA that facilitates scalable LoRA fusion within LLMs, enhancing the existing LLMs to be an MoE-style architecture that fuses multiple LoRA adapters as experts with an efficient operator to accelerate the forward process. LLMs equipped with MeteoRA enhance the ability to automatically discern the task intention from inputs and select the most pertinent LoRA adapter to generate appropriate responses. Moreover, its capability for timely intention switching leads to superior performance, particularly in sequentially solving composite tasks. Future work could explore the transformative potential of MeteoRA in multifaceted problem-solving scenarios, especially those requiring the fusion of diverse knowledge from individually fine-tuned LoRA adapters.
@@ Appendix A Appendix
A.1 28 tasks details
@@
| abstract_narrative_understanding | narrative understanding, multiple choice | Given a narrative, choose the most related proverb. | Accuracy |
@@
| alpaca | instruction-tuning | Write appropriate answers according to instructions. | BLEU, ROUGE |
@@
| cnn_dailymail | summarization | Given news articles, write the summarization. | ROUGE |
@@
| contextual_parametric_knowledge_conflicts | contextual question-answering, multiple choice | Answer questions given the contextual information. | Accuracy |
@@
| cs_algorithms | algorithms, numerical response | Solve two common computer-science tasks. | Accuracy |
@@
| disfl_qa | contextual question-answering, reading comprehension | Pick the correct answer span from the context given the disfluent question. | Accuracy |
@@
| elementary_math_qa | mathematics | Answer multiple choice mathematical word problems. | Accuracy |
@@
| epistemic_reasoning | logical reasoning, multiple choice | Determine whether one sentence entails the next. | Accuracy |
@@
| formal_fallacies_syllogisms_negation | logical reasoning, multiple choice, | Distinguish deductively valid arguments from formal fallacies. | Accuracy |
@@
| goal_step_wikihow | causal reasoning, multiple choice | Perform one of three subtasks: step inference, goal inference, or step ordering. | Accuracy |
@@
| gsm8k | mathematics | Solve the grade school math word problems. | Accuracy |
@@
| language_identification | multilingual, multiple choice | Given a sentence, select the correct language. | Accuracy |
@@
| linguistics_puzzles | logical reasoning, linguistics | Solve Rosetta Stone-style linguistics puzzles. | BLEU, ROUGE |
@@
| logical_deduction | logical reasoning, multiple choice | Deduce the order of a sequence of objects. | Accuracy |
@@
| news_commentary_de | multilingual, translation | Translate German sentences into English. | BLEU |
@@
| news_commentary_es | multilingual, translation | Translate Spanish sentences into English. | BLEU |
@@
| news_commentary_it | multilingual, translation | Translate Italian sentences into English. | BLEU |
@@
| object_counting | logical reasoning | Questions that involve enumerating objects and asking the model to count them. | Accuracy |
@@
| paragraph_segmentation | segmentation, multilingual | Identify the sentences that end a paragraph in a document. | Accuracy |
@@
| play_dialog_same_or_different | reading comprehension, multiple choice | Determine if nearby lines in a Shakespeare play were spoken by the same individual. | Accuracy |
@@
| question_selection | reading comprehension, multiple choice | Given an answer along with its context, select the most appropriate question which has the given answer as its answer. | Accuracy |
@@
| reasoning_about_colored_objects | reading comprehension, logical reasoning, multiple choice | Answer extremely simple questions about the colors of objects on a surface. | Accuracy |
@@
| strategyqa | logical reasoning, context-free question answering | Answer questions in which the required reasoning steps are implicit in the question. | BLEU, ROUGE, Accuracy |
@@
| topical_chat | free response | Open-domain response generation. | BLEU, ROUGE |
@@
| tracking_shuffled_objects | logical reasoning, multiple choice | Determine the final positions given initial positions and a description of a sequence of swaps. | Accuracy |
@@
| unit_conversion | contextual question-answering, mathematics, multiple choice | Perform various tasks relating to units, including identification and conversion. | Accuracy |
@@
| vitaminc_fact_verification | truthfulness, reading comprehension, multiple choice | Identify whether a claim is True or False based on the given context. | Accuracy |
@@
| winowhy | causal reasoning, multiple choice | Evaluate the reasoning in answering Winograd Schema Challenge questions. | Accuracy |
A.2 28 tasks evaluation result details

### v2

@@ 4.2 Main results
Figures 3(a) and 3(b) demonstrate the performance of the MeteoRA models, the compared methods (LoRA-F and LoRA-B), and a reference model PEFT based on LlaMA2-13B and LlaMA3-8B, respectively, across the selected 28 tasks. Table 1 shows the averaged scores in various matrics for all methods.
@@
| PEFT (reference) | 0.762 / 0.817 | 35.66 / 45.32 | 0.340 / 0.341 | 0.163 / 0.164 | 0.316 / 0.317 |
| MeteoRA (top-1) | 0.748 / 0.806 | 41.74 / 44.98 | 0.336 / 0.337 | 0.160 / 0.158 | 0.314 / 0.313 |
| MeteoRA (top-2) | 0.754 / 0.811 | 36.73 / 45.64 | 0.336 / 0.338 | 0.160 / 0.158 | 0.313 / 0.314 |
| LoRA-F | 0.729 / 0.767 | 41.27 / 42.93 | 0.318 / 0.327 | 0.136 / 0.157 | 0.294 / 0.306 |
| LoRA-B | 0.665 / 0.750 | 37.98 / 38.47 | 0.318 / 0.343 | 0.128 / 0.171 | 0.288 / 0.320 |
The evaluation results indicate that, regardless of the base LLM, the MeteoRA models utilizing the top-1 strategy achieve performance very close to the reference model PEFT, while no explicit LoRA activation/deactivation is required in MeteoRA. Although LLMs with both LoRA-F and LoRA-B reach comparable performance on several certain tasks, they exhibit significantly poorer outcomes on others. Additionally, MeteoRA employing the top-2 strategy, despite occasionally showing greater capability loss compared to MeteoRA with top-1 strategy, occasionally outperforms PEFT with adapters trained directly on the individual tasks. This suggests that the component in the loss function (Equation 4) becomes influential in these cases, indicating a beneficial mix of LoRA adapters from various tasks for future study.
@@ 4.3 Composte-n tasks
The evaluation results for these three tasks are illustrated in Table 2. Notice that only LlaMA3-8B with the MeteoRA (top-2 strategy) and LoRA-B effectively address these composite-n tasks. Subsequent discussions will therefore focus exclusively on these two models. Although the MeteoRA model attempts slightly fewer questions than LoRA-B in composite-3 tasks, it correctly answers a higher number of multiple-choice questions and achieves superior BLEU and ROUGE scores. As the task complexity increases to composite-5 and composite-10, MeteoRA outperforms LoRA-B in almost all metrics. For more details, refer to Appendix A.5.
@@
To further validate the functionality of the Gating network in the MeteoRA block, we display the LoRA selection patterns in the inference process of a composite-3 sample in Figure 4. With the top- strategy, Gating network appropriately assigns greater weight to the corresponding LoRA adapters for the majority of the tokens, no matter in input or output. At the junctions of two adjacent tasks, the Gating network correctly performed the timely switching actions of LoRA adapters.
4.4 Efficiency
To assess the efficiency of our novel forward pass designs using custom GPU kernel operators, we truncate samples from each test dataset of all 28 tasks. We evaluate these designs alongside four variants with the same hyperparameters: the upper-bound single-lora, the baseline loop-original, and two novel forward acceleration strategies based on : bmm-torch and bmm-triton, implemented by PyTorch and Triton respectively. Figure 5 displays the histogram of the overall root-of-runtime metric for each task and design. Additional efficiency evaluation experiments are detailed in Appendix A.7.
@@ 5 Related Work
Multi-task fusion. Our proposed method falls into the field of LoRA adapter composition for multi-task fusion. The first category focuses on fusing the entire models. Researchers mainly study model ensembling and multi-task learning to achieve this goal. Existing works integrate the models under the setting of shared model architecture (Matena and Raffel, 2022; Jin et al., 2022; Wu et al., 2023c; Yadav et al., 2024). Others focus on merging models with various architectures or from different tasks. Both methods (Stoica et al., 2023; Liu et al., 2022) try to merge models that are trained for various tasks without additional training. The second category is more concerned with fusion in terms of the tasks. Ilharco et al. (2022) proposes a model editing method via task vectors. Sun et al. (2022) leverage in-context learning with few-shots to enhance the performance of unseen tasks. However, these methods require multi-task training or prior knowledge for the evaluation tasks. Our method embeds off-the-shelf LoRA adapters with a Gate network in the MeteoRA module. None of the examples (zero-shot) are required for all individual tasks.
@@
Fusion under MoE. In the context of pretrain-fine-tune paradigm, PEFT becomes a common sense for developing Transformer-based LLM downstream applications. Directly fine-tuning on a fused dataset from various tasks is unable to achieve better performance Ling et al. (2024). Some works focus on leveraging existing LoRA adapters as off-the-shelf components, integrating them directly into a base LLM. For example, PEFT (Mangrulkar et al., 2022) and S-LoRA (Sheng et al., 2023) are frameworks aiming to embed multiple LoRA adapters to one LLM. However, requiring explicit activation/deactivition during usage. MixLoRA (Li et al., 2024) targets to a resource-efficient sparse MoE model, fine-tuning MoE on MLP module with the auxiliary load balance loss used in Mixtral (Jiang et al., 2024). Although MixLoRA supports LoRA adapters for the attention layer, the adapters are still dense models encompassed with the linear layers in the attention module. Others (Huang et al., 2023; Yang et al., 2024; Feng et al., 2024; Chen et al., 2024; Wu et al., 2023d) propose LoRA fusion based on the concept of Mixture-of-experts that enhance the model’s ability for cross-domain tasks. However, the methods mainly focus on fusing LoRA adapters to the FFN module or Q in the attention module. Our method could embed all kinds of LoRA adapters. By leveraging the full-mode MoE architecture, the LLM’s capacity could be boosted with autonomous and timely LoRA switching, especially for solving composite tasks.
6 Limitations
LoRA adapter update. Although the Gating network within MeteoRA module is trained separately among the adapters, it is necessary to retrain or fine-tune the Gating network if some LoRA adapters are updated. The Gating network is trained using the hidden state as inputs, which are influenced by LoRA adapters in previous layers. Testing revealed that directly replacing some LoRA adapters with improved versions did not enhance performance on our test set. However, after retraining the MeteoRA modules, the LLM equipped with MeteoRA exhibited performance improvements. Technically, this issue may be related to the domain shift problem, where the Gating network is applied to another operational field the distribution shift. Employing statistical methods such as (Li et al., 2020; Krishnan and Tickoo, 2020) may help calibrate the output of the Gating network to produce more accurate logits and results.
Knowledge fusion tasks. Composite tasks, which involve a broad range of tasks, represent one type of complexity in terms of the scope of tasks. More challenging are tasks that require knowledge fusion across domains. To assess the capability of MeteoRA in knowledge fusion task, we construct a mathematics task by translating problems from GSM8K into a foreign language (e.g., Italian), so that the LLM with MeteoRA must solve these foreign language GSM8K problems by leveraging knowledge from both GSM8k LoRA (trained on problems in English) and the foreign language LoRA (trained for Italian to English translation). Although MeteoRA successfully fuses the two LoRA adapters to address the math problems in a foreign language, it does not show superior performance compared to LLM equipped only with the GSM8K LoRA. We hypothesize that the base LLM’s existing proficiency in the selected foreign language may render the additional adapter unnecessary. Future efforts could focus on constructing more suitably complex tasks where the required cross-domain knowledge is not already pre-trained into the base LLM.
MoE efficiency. Sparsely-gated MoE (Shazeer et al., 2017) offers computational efficiency advantages over dense MoE. However, the naive implementation of MoE forward (loop-original), such as the SparseMoE in Mixtral (Jiang et al., 2024; Wolf et al., 2020), still encounters efficiency issues when the number of experts increases. In our evaluations, the runtime for inference can be up to longer than that of single-lora when embedding 28 LoRA adapters into one LLM. With our proposed forward acceleration techniques bmm-torch and bmm-triton, we achieve a speedup of compared to the loop-original, though this still falls short of the ideal upper bound (single-lora). Technically, it is extremely difficult to increase the inference speeds for MeteoRA when the number of embedded LoRA adapters increases. Future work could explore developing new operators in triton or CUDA to continuously enhance MoE acceleration in terms of memory efficiency.
@@ 7 Conclusions
This paper presents a framework MeteoRA that achieves scalable multi-task LoRA embedding within LLMs, enhancing the existing LLMs with a full-mode MoE architecture with forward acceleration strategies. LLMs equipped with MeteoRA enhance the ability to autonomously select the most pertinent LoRA adapters to generate appropriate responses. Moreover, its capability for timely LoRA switching leads to superior performance, particularly in sequentially solving composite tasks. Future work could explore the transformative potential of MeteoRA in multifaceted problem-solving scenarios, and inference efficiency by designing more efficient GPU kernel operators.
@@ Appendix A Appendix
A.1 Top-k strategy
Top- strategy: With the top- strategy set in the Gating network, MeteoRA computes the normalized weights for the selected LoRA adapters. These weights participate in the computation of the losses of both and as specified in Equation 2. Thus, the parameter updates for the Gating network derive from the losses associated with both LoRA classification and autoregressive token generation. Notice that the LoRA classification loss is only influenced by the LoRA adapter with the highest logit, whereas the backpropagation from the token loss affects the parameters in the Gating network responsible for all selected LoRA adapters. Although these remaining adapters lack direct supervision from the LoRA classification loss, the token generation loss contributes to enhanced robustness and the capacity for LoRA switching during generation.
A.2 Details on the tailored Triton kernel for efficient MeteoRA forward
To address the memory copying problem caused by PyTorch indexing, we fuse the two operations inside a GPU kernel function implemented by Triton, which dynamically indexes the right pair of LoRA matrix and load them from HBM to SRAM in each parallelized thread. Therefore, there is no need to explicitly allocate pairs of over the original ones.
Another challenge is that Triton constraints all the dimensions for the matrix operators should be no less than , however, under the MeteoRA settings, this requirement can never be satisfied since the first operator is a vector, and also, the LoRA rank size may be less than easily (e.g., in all our experiments, we fix ). Therefore, it is not that trivial to implement such a kernel, unless using the simple masking strategy to meet the requirements with over waste of I/O.
To both obey the dimension constraint and avoid too much waste, for the first of and , we use a blocking strategy to split the vector along the hidden size dimension by blocks, where . In such case, the first operator becomes a matrix with shape , and also we have to split along the first dimension to become a more square matrix with shape . Notice that now the output of first : with shape has a relationship with the original one with shape as follows:
where and are two trivial mask matrixs, each sized and respectively. So we can transform back to the right results by three additional negligible dot-product with , matrix-product with , and colsum operations for the first . For the second one, instead of directly using the right result , we can delay the colsum operation until we finish the second , i.e. we use with shape and with shape to do matrix-product operations to get the temporary result with shape , then apply colsum to get the final LoRA output with shape . Notably, on one hand, we can avoid one more blocking operation for since already meets the dimension constraint, on the other hand, if , we can just simply utilize masking strategy since it is the inner dimension and small enough.
Overall, for the Triton kernel function, we offer the pseudo code as shown in Algorithm 1.
A.3 Information about 28 tasks
Table 3 shows the detailed information of the 28 selected tasks in the Section 4. The name in parentheses is the abbreviation of the corresponding task. We use the original training sets from these tasks to fine-tune the LoRA adapters and Gating networks in MeteoRA modules. To achieve a balanced fine-tuning across the diverse task spectrum and ensure efficient training, we construct a balanced dataset by randomly sampling 1,000 samples from each task. This balanced dataset is then divided into a training set with 25,200 samples (i.e., 900 samples for each task) and a validating set with 2,800 samples (i.e., 100 samples for each task) for fine-tuning. In terms of the evaluation part, the performances are evaluated on each task’s original test set.
@@ A.3 Information about 28 tasks
| abstract_narrative_understanding (AbsNarr) | narrative understanding, multiple choice | Given a narrative, choose the most related proverb. | Accuracy |
@@
| alpaca (ALPACA) | instruction-tuning | Write appropriate answers according to instructions. | BLEU, ROUGE |
@@
| cnn_dailymail (CNNDM) | summarization | Given news articles, write the summarization. | ROUGE |
@@
| contextual_parametric_knowledge_conflicts (ConParaKC) | contextual question-answering, multiple choice | Answer questions given the contextual information. | Accuracy |
@@
| cs_algorithms (CSAlg) | algorithms, numerical response | Solve two common computer-science tasks. | Accuracy |
@@
| disfl_qa (DisflQA) | contextual question-answering, reading comprehension | Pick the correct answer span from the context given the disfluent question. | Accuracy |
@@
| elementary_math_qa (ElemMath) | mathematics | Answer multiple choice mathematical word problems. | Accuracy |
@@
| epistemic_reasoning (EpiReason) | logical reasoning, multiple choice | Determine whether one sentence entails the next. | Accuracy |
@@
| formal_fallacies_syllogisms_negation (FormFall) | logical reasoning, multiple choice, | Distinguish deductively valid arguments from formal fallacies. | Accuracy |
@@
| goal_step_wikihow (GSWiki) | causal reasoning, multiple choice | Perform one of three subtasks: step inference, goal inference, or step ordering. | Accuracy |
@@
| gsm8k (GSM8K) | mathematics | Solve the grade school math word problems. | Accuracy |
@@
| language_identification (LangID) | multilingual, multiple choice | Given a sentence, select the correct language. | Accuracy |
@@
| linguistics_puzzles (LingPuzz) | logical reasoning, linguistics | Solve Rosetta Stone-style linguistics puzzles. | BLEU, ROUGE |
@@
| logical_deduction (LogDeduc) | logical reasoning, multiple choice | Deduce the order of a sequence of objects. | Accuracy |
@@
| news_commentary_de (NewsDE) | multilingual, translation | Translate German sentences into English. | BLEU |
@@
| news_commentary_es (NewsES) | multilingual, translation | Translate Spanish sentences into English. | BLEU |
@@
| news_commentary_it (NewsIT) | multilingual, translation | Translate Italian sentences into English. | BLEU |
@@
| object_counting (ObjCount) | logical reasoning | Questions that involve enumerating objects and asking the model to count them. | Accuracy |
@@
| paragraph_segmentation (ParaSeg) | segmentation, multilingual | Identify the sentences that end a paragraph in a document. | Accuracy |
@@
| play_dialog_same_or_different (PlayDiag) | reading comprehension, multiple choice | Determine if nearby lines in a Shakespeare play were spoken by the same individual. | Accuracy |
@@
| question_selection (QuestSel) | reading comprehension, multiple choice | Given an answer along with its context, select the most appropriate question which has the given answer as its answer. | Accuracy |
@@
| reasoning_about_colored_objects (ColorReason) | reading comprehension, logical reasoning, multiple choice | Answer extremely simple questions about the colors of objects on a surface. | Accuracy |
@@
| strategyqa (StratQA) | logical reasoning, context-free question answering | Answer questions in which the required reasoning steps are implicit in the question. | BLEU, ROUGE, Accuracy |
@@
| topical_chat (TopChat) | free response | Open-domain response generation. | BLEU, ROUGE |
@@
| tracking_shuffled_objects (TrackObj) | logical reasoning, multiple choice | Determine the final positions given initial positions and a description of a sequence of swaps. | Accuracy |
@@
| unit_conversion (UnitConv) | contextual question-answering, mathematics, multiple choice | Perform various tasks relating to units, including identification and conversion. | Accuracy |
@@
| vitaminc_fact_verification (VitaFact) | truthfulness, reading comprehension, multiple choice | Identify whether a claim is True or False based on the given context. | Accuracy |
@@
| winowhy (WinoWhy) | causal reasoning, multiple choice | Evaluate the reasoning in answering Winograd Schema Challenge questions. | Accuracy |
A.4 Experimental results of 28 tasks